# Anomalous strain rate sensitivity of a Cu/Al2O3 multi-layered thin film

### v1

@@ 4.1 Microstructure evolution
As it can be seen in Fig. 4,and Suppl. Figs. S5 and S7, the heat treatment has increased the initial grain size, that has caused the previously flat Al2O3 layers to become rippled, following the curvature of the grains growing in their vicinity. This process can also affect the thickness of the Al2O3 layer, too, as the increased grain surface can thin down the interlayer, just like the material of a balloon upon its inflation. One of the striking things to notice was that the grain growth has seemingly reached a final state and further grain growth was blocked even at longer ageing (72 hours was the longest achieved). Even after repeated attempts (initially at 600C, then 700C and finally 800C treatment using different testpieces), the grain growth process was halted. In this system, the Cu grain boundaries could not disappear due to an unknown phenomena. As a possible explanation, the diffusion of the Al (or Al2O3) has been proposed, by either i) precipitation along grain boundaries, or ii) atomic diffusion, as in an intra- or inter-granular manner. High resolution TEM imaging has confirmed the first hypothesis, showing that Al2O3 precipitates are located between large grains (marked by white arrows in Fig. 4c), where the curvature of the oxide layer has significantly increased. In order to confirm the second hypothesis based on atomic level diffusion, further experiments will have to be performed, that is out of the scope of the current work. In summary, the aluminum-oxide layer has successfully served its purpose as a thickness (and grain size) controlling barrier, but if larger grains are needed to be designed in the future, the composition or the thickness of the interlayer will need to be modified in the currently investigated system.
@@ 4.2 Strain rate related microstructure transformation
In Ref. [18], when UFG pillars were compressed at HSR, some pillars exhibited grain coarsening in the gauge section, that was attributed to static recrystallization due to the extremely high dislocation densities, and not resulting from dynamical processes during the compression itself [37, 38]. In order to verify the appearance of recrystallization in this thin film (that could potentially modify the deposited coating’s properties after impact), the deformed pillars were mapped by EBSD immediately ($\sim 2$ hours) after micromechanical testing (Suppl. Figs. S5- S7), and after ageing the samples under ambient conditions for 1.5 years (Fig. S8). We found no evidence of abnormal grain growth in these pillars, regardless of the applied SR. The absence of recrystallization possibly originates from the applied Al2O3 interlayers, that conclusively not only exhibit excellent high temperature stability, but could potentially deactivate unwanted microstructure transformation, if the coating is to be operating under extreme conditions.
@@ 5 Conclusions
During the mechanical data analysis, two anomalous phenomena was identified in the strain rate sensitivity behavior, namely i) yield stress saturation in the AD samples when the pillar diameters were below a given threshold, and ii) increased strain rate hardening in the HT samples at very high ($\dot{\varepsilon}=1000$/s) strain rates. This behavior clearly demonstrates that dislocation nucleation and motion is strongly deformation rate dependent in Cu at high strain rates.
@@
GND densities were also studied in detail before and after compression, leading to the observation of slightly higher dislocation densities in the vicinity of the pillars’ edges, resulting in an inhomogeneous distribution within the samples. This most likely plays a key role in the mechanical behavior and results in a transition of the strain rate sensitivity from monotonous to anomalous, when the sample sizes are reduced below a characteristic (microstructure dependent) length scale.
@@ Acknowledgement
Anomalous strain rate sensitivity of a Cu/Al2O3 multi-layered thin film
Szilvia Kalácskaa,b,∗, László Pethőb, Guillaume Kermouchea, Johann Michlerb, Péter D. Ispánovityc
@@
b Empa, Swiss Federal Laboratories for Materials Science and Technology, Laboratory of Mechanics of Materials and Nanostructures, CH-3602 Thun, Feuerwerkerstrasse 39. Switzerland
@@
Suppl. Figure S 4: Yield stress determination procedure, pillar ”AD” P01. This method was used to be consistent even at high strain rates, where the ”ringing effect” introduces oscillations to the stress-strain curve.
Suppl. Figure S 5: EBSD results, 4 $\upmu$m (big) pillars. Z directional inverse pole figure (IPF Z) colouring is shown together with band contract (BC, greyscale) imaging, before and after deformation. In the to right corner the full IPF triange is shown along with the registered grain colours plotted inside the IPF triangle for the as deposited (AD) case. Pole figures plotted for the X, Y and D directions show no specific texture in the AD sample. Red dots on the EBSD maps highlight the two pillars that were analysed by post mortem 3D HR-EBSD.
Suppl. Figure S 6: EBSD results, 4.0 $\upmu$m pillars, 2nd batch (for more statistics). Z directional inverse pole figure (IPF Z) colouring is shown together with band contract (BC, greyscale) imaging, before and after deformation. ”AA crash” refers to the accidental touch of the pillar upon auto approach procedure, which resulted in an unknown deformation state, therefore this pillar was excluded from the analysis.
Suppl. Figure S 7: EBSD results, 2.9 $\upmu$m pillars (for size effect). Z directional inverse pole figure (IPF Z) colouring is shown together with band contract (BC, greyscale) imaging, before and after deformation.
Suppl. Figure S 8: EBSD results of the deformed AD pillars, aged for 1.5 years, repolished by FIB.
Suppl. Figure S 9: TEM lamella preparation. a) snapshot during the liftout process, b) image taken during the thinning.
Suppl. Figure S 10: Graphical Abstract

### v2

@@ 4.1 Microstructure evolution
As it can be seen in Fig. 4,and Suppl. Figs. S6 and S8, the heat treatment has increased the initial grain size, that has caused the previously flat Al2O3 layers to become rippled, following the curvature of the grains growing in their vicinity. This process can also affect the thickness of the Al2O3 layer, too, as the increased grain surface can thin down the interlayer, just like the material of a balloon upon its inflation. One of the striking things to notice was that the grain growth has seemingly reached a final state and further grain growth was blocked even at longer ageing (72 hours was the longest achieved). Even after repeated attempts (initially at 600C, then 700C and finally 800C treatment using different testpieces), the grain growth process was halted. In this system, the Cu grain boundaries could not disappear due to an unknown phenomena. As a possible explanation, the diffusion of the Al (or Al2O3) has been proposed, by either i) precipitation along grain boundaries, or ii) atomic diffusion, as in an intra- or inter-granular manner. High resolution TEM imaging has confirmed the first hypothesis, showing that Al2O3 precipitates are located between large grains (marked by white arrows in Fig. 4c), where the curvature of the oxide layer has significantly increased. In order to confirm the second hypothesis based on atomic level diffusion, further experiments will have to be performed, that is out of the scope of the current work. In summary, the aluminum-oxide layer has successfully served its purpose as a thickness (and grain size) controlling barrier, but if larger grains are needed to be designed in the future, the composition or the thickness of the interlayer will need to be modified in the currently investigated system. Regarding the role of the Cu/alumina interface, further high resolution studies are needed to understand the extent to which the crystalline-amorphous interface in a metal-ceramic hybrid system influences dislocation nucleation, mobility, and annihilation.
@@ 4.2 Strain rate related microstructure transformation
In Ref. [18], when UFG pillars were compressed at HSR, some pillars exhibited grain coarsening in the gauge section, that was attributed to static recrystallization due to the extremely high dislocation densities, and not resulting from dynamical processes during the compression itself [46, 47]. In order to verify the appearance of recrystallization in this thin film (that could potentially modify the deposited coating’s properties after impact), the deformed pillars were mapped by EBSD immediately ($\sim 2$ hours) after micromechanical testing (Suppl. Figs. S6- S8), and after ageing the samples under ambient conditions for 1.5 years (Fig. S9). We found no evidence of abnormal grain growth in these pillars, regardless of the applied SR. The absence of recrystallization possibly originates from the applied Al2O3 interlayers, that conclusively not only exhibit excellent high temperature stability, but could potentially deactivate unwanted microstructure transformation, if the coating is to be operating under extreme conditions.
@@ 5 Conclusions
During the mechanical data analysis, the strain rate sensitivity behavior was analyzed in detail, namely i) yield stress as a function of strain rate and pillar diameters, and ii) increased strain rate hardening in the HT samples at very high ($\dot{\varepsilon}=1000$/s) strain rates. It was proposed that this behavior is a demonstration that the dislocation nucleation and motion is strongly deformation rate dependent in Cu at high strain rates.
@@
GND densities were also studied in detail before and after compression, leading to the observation of slightly higher dislocation densities in the vicinity of the pillars’ edges, resulting in an inhomogeneous distribution within the samples. This most likely plays a key role in the mechanical behavior and results in a transition of the strain rate sensitivity at HSR, when the sample sizes are reduced below a characteristic (microstructure dependent) length scale.
@@ Acknowledgement
Strain rate sensitivity of a Cu/Al2O3 multi-layered thin film
Szilvia Kalácskaa,b,∗, Laszlo Pethöb, Guillaume Kermouchea, Johann Michlerb, Péter D. Ispánovityc
@@
b Empa, Swiss Federal Laboratories for Materials Science and Technology, Laboratory for Mechanics of Materials and Nanostructures, Feuerwerkerstrasse 39, 3602 Thun, Switzerland
@@
a)Refer to captionb)
Suppl. Figure S 4: a) Yield stress determination procedure, pillar “AD” P01. This method was used to be consistent even at high strain rates, where the ”ringing effect” introduces oscillations to the stress-strain curve. b) Figure 2a engineering stress – engineering strain values converted to $\sigma_{True}-\varepsilon_{True}$ curves. The applied compliance correction is $displacement-load\times C_{i}$, where $C_{STD}=3.695\times 10^{-6}$, $C_{MLC}=8\times 10^{-7}$ and $C_{ST}=9.5\times 10^{-7}$.
a)Refer to captionb)
c)Refer to caption
Suppl. Figure S 5: a) Comparison of the individual yield stresses determined by the linear fitting method (empty signs, used in Fig. 3) and the conventional 1% plastic strain offset method (filled signs). b) (Same as a), but here the average values were plotted (same as Fig.3a), excluding the 1000/s results. c) ) (Same as Fig.3b). The 1% method results in a similar m value for the AD case ($m_{\mathrm{AD}}=0.02\pm 0.01)$, however, the HT evaluation becomes much more noisy, resulting in a negative $m_{\mathrm{HT}}$ value ($m_{\mathrm{HT}}=-0.003)$, that is physically not expected and originating from the increased error in the determined $\sigma_{y}$.
Suppl. Figure S 6: EBSD results, 4 $\upmu$m (big) pillars. Z directional inverse pole figure (IPF Z) colouring is shown together with band contract (BC, greyscale) imaging, before and after deformation. In the to right corner the full IPF triange is shown along with the registered grain colours plotted inside the IPF triangle for the as deposited (AD) case. Pole figures plotted for the X, Y and D directions show no specific texture in the AD sample. Red dots on the EBSD maps highlight the two pillars that were analysed by post mortem 3D HR-EBSD.
Suppl. Figure S 7: EBSD results, 4.0 $\upmu$m pillars, 2nd batch (for more statistics). Z directional inverse pole figure (IPF Z) colouring is shown together with band contract (BC, greyscale) imaging, before and after deformation. ”AA crash” refers to the accidental touch of the pillar upon auto approach procedure, which resulted in an unknown deformation state, therefore this pillar was excluded from the analysis.
Suppl. Figure S 8: EBSD results, 2.9 $\upmu$m pillars (for size effect). Z directional inverse pole figure (IPF Z) colouring is shown together with band contract (BC, greyscale) imaging, before and after deformation.
Suppl. Figure S 9: EBSD results of the deformed AD pillars, aged for 1.5 years, repolished by FIB.
Suppl. Figure S 10: TEM lamella preparation. a) snapshot during the liftout process, b) image taken during the thinning.
Suppl. Figure S 11: Graphical Abstract